I need to use the number 7 four times as exponents to equal 1, then 2, then 3...all the way up to 100. How do I figure out the equations?
For example: 77/77 = 1
7/7 + 7/7 = 2 and so on...

Answer by rapaljer(4671)

This sounds like fun! Get as many different combinations as you can think of, and see which ones are missing. Here are a that I could think of:

77/7-7=4
(7-7)/7 +7 = 7
((7*7+7)/7 =8
(7+7)/7 +7 =9
(77-7)/7 =10
(77+7)/7 =12
(7*7)/7+7=14
77/7+7=18
7+7+7+7 =28
7*7 +7/7 = 50
77-7/7 = 76
77+7/7=78
77+7+7=91
This sounds like a good group project or a contest to see who can come up with the most different numbers. Am I right??

Have fun with the rest of these!!

R^2

Dr. Robert J. Rapalje, Retired
Seminole Community College
Altamonte Springs Campus
Florida